OpenStudy (anonymous):

I need some help with checking this answer please: \[\large\int\limits_4^5 \frac{x^3-3 x^2-9}{x^3-3 x^2} dx = x-\frac{3}{x}-\ln|3-x|+\ln|x|+C\] I got this far using the partial fractions technique. I went to check my answer that I got on paper but Wolfram gives some craziness with hyperbolic tangents that I don't understand. So what should the answer be and wouldn't the natural log of 3-x be an issue because it gives non-real values? The correct answer MUST have real values and be in terms of constants and logs. If somebody could show me what's going on here it would be great :-)

OpenStudy (anonymous):

For my partial fractions I got: \[x^3-3x^2-9=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x-3}\] The A part turns into ln|x|, the B part turns into 3/x, the C part turns into -ln|3-x|

OpenStudy (helder_edwin):

\[ \Large \int\frac{x^3-3x^2-9}{x^3-3x^2}\,dx= \int\,dx-\int\frac{9}{x^2(x-3)}\,dx \] so you have to solve \[ \Large \frac{9}{x^2(x-3)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x-3} \]

OpenStudy (valpey):

Well, for one thing you have written the natural log of the absolute value of 3-x.

OpenStudy (anonymous):

Yeah I used long division initially to pull out this: \[1 + \frac{-9}{x^3-3x^2}\]

OpenStudy (anonymous):

Integral of one with respect to x is just x, logically

OpenStudy (anonymous):

I think I mistyped my second post, sry @helder_edwin

OpenStudy (anonymous):

@Valpey oh yeah... so it's always a positive...

OpenStudy (anonymous):

ln|3-(4)| = ln|-1| = ln 1 = 0

OpenStudy (anonymous):

That's a nice simplification

OpenStudy (anonymous):

/me waves hello @asnaseer :-) ty for stopping by!

OpenStudy (asnaseer):

you didn't need to do long division to simplify this in the first place. you could just have done this:\[\frac{x^3-3x^2-9}{x^3-3x^2}=\frac{x^3-3x^2}{x^3-3x^2}-\frac{9}{x^3-3x^2}=1-\frac{9}{x^3-3x^2}\]

OpenStudy (asnaseer):

secondly, the logs should not contain the absolute value symbols around them

OpenStudy (anonymous):

Good point @asnaseer and why not? Without it wouldn't I have a non-real answer issue when I go to evaluate?

OpenStudy (asnaseer):

you should have done this:\[\int\limits_4^5 \frac{x^3-3 x^2-9}{x^3-3 x^2} dx =[ x-\frac{3}{x}-\ln(3-x)+\ln(x)]_4^5\]\[\qquad=[x-\frac{3}{x}+ln(\frac{x}{3-x})]_4^5\]\[\qquad=(5-\frac{3}{5}+ln(-\frac{5}{2}))-(4-\frac{3}{4}+ln(-\frac{4}{1}))\]\[\qquad=5-\frac{3}{5}-4+\frac{3}{4}+ln(-\frac{5}{2}\div-\frac{4}{1})\]

OpenStudy (asnaseer):

and the minuses cancel out inside the logs

OpenStudy (anonymous):

*reading carefully*

OpenStudy (asnaseer):

BTW: waves hello back to @agentx5 :)

OpenStudy (anonymous):

Oh... properties of logs... \[\ln|a|-\ln|b|=\ln|\frac{a}{b}|\]

OpenStudy (asnaseer):

yes - but you shouldn't use the '|' symbols here

OpenStudy (asnaseer):

as they usually indicate absolute value

OpenStudy (anonymous):

\[\qquad=5-\frac{3}{5}-4+\frac{3}{4}+ln(-\frac{5}{2}\div-\frac{4}{1}) \approx 3.45258\] Disagrees with this check: http://www.wolframalpha.com/input/?i=integrate+from+4+to+5+for+%28x^3-3x^2-9%29%2F%28x^3-3x^2%29 \(\approx \frac{17}{25}\)

OpenStudy (asnaseer):

I work it out to be exactly the same as the answer wolfram gives

OpenStudy (asnaseer):

I get:\[\frac{23}{20}+ln(\frac{5}{8})\approx 0.68\]

OpenStudy (anonymous):

Ugh I see now I used the wrong base log on my calculator... /facepalm

OpenStudy (anonymous):

Derp.

OpenStudy (asnaseer):

np :)

OpenStudy (anonymous):

Thank you so much you were a great help! @TuringTest give that claymation-man a medal ^_^

OpenStudy (asnaseer):

you are more than welcome my friend! :)